OpenStudy (anonymous):

3^x=e^2 do i think ln both side first?

hartnn (hartnn):

@elskydiablo Hi, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) taking ln on both sides is a good idea, , you'll just get 2 on the right. But what exactly you want to do ? solve for 'x' ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

great, so can you simplify ln (3^x) ?? use \(\large \ln a^b = b\ln a\)

OpenStudy (anonymous):

x ln(3)=2 and i just divide ln(3) both side and that my answer right?

hartnn (hartnn):

absolutely correct! good :)

OpenStudy (anonymous):

thank you so much for helping

hartnn (hartnn):

you are most welcome ^_^